OpenStudy (anonymous):

Can someone help me with these two problems please. Create your own binomial expression with a radical in the second term. Question 1: Identify its conjugate and explain, in complete sentences, why it is the conjugate. Question 2: Multiply your original binomial expression and its conjugate. What happened to the radicals and why?

OpenStudy (anonymous):

So you just want to make up some irrational number. Say its: \[3+\sqrt{2}\] Its conjugate would be: \[3-\sqrt{2}\] When you multiply them (using the FOIL technique) you get: \[(3+\sqrt{2})(3-\sqrt{2}) = (3)(3)-(3)(\sqrt{2})+(3)(\sqrt{2})-(\sqrt{2})(\sqrt{2})\] Noitice that the two middle terms will cancel out, and the last term wont be irrational anymore (this always happens when multiplying an irrational number with its conjugate)